Weekly Science Report 6: LMG 02-05 Southern Ocean GLOBEC

Cruise days 2-9 September inclusive


BG-232-0 Costa/Burns/Crocker – Foraging Ecology of Crabeater seals

BG-234-0 Fraser – Winter Foraging Ecology of Adélie Penguins

BG-235-0 Fritsen – Sea Ice Microbial Communities

OG-241-0 Smith/Martinson/Perovich – Optical Environment of the Western Antarctic Peninsula Region

BG-244-0 Quetin/Ross – Winter Ecology of Larval Krill: Quantifying their interaction with the Pack Ice habitat


Cruise overview to date:


The week began abruptly with a minor incident involving the Dush-5 winch, though no one was injured and only minimal equipment damage was incurred.  The problem was quickly traced to a faulty winch control, which was replaced and the CTD system checked over before the evening cast using the Dush-4/fantail CTD configuration.  The remainder of the week was rather unremarkable but very productive despite our needing to relocate the LMG several times when our floes broke apart or the ship could not hold position with light winds and loose pack ice.  All science groups were able to conduct operations including the Seal team, which had not seen many workable animals in some time.  We are beginning to wrap up operations, pack gear, and prepare for our port calls at Palmer Station and Punta Arenas.  Meanwhile, an Antarctic Circle Crossing Ceremony is in the works.


Daily Activities:


2 September 2002

The LMG continues to occupy Process Station #3.  This site is unlike our previous two locations because we are able to move about on our own.  There are no sufficiently large ice floes upon which we can maintain a continued presence.  Therefore, we have been picking a “floe of the day” each morning and gathering as much data as possible during the day before moving off to a new floe the following day.  Also unlike our previous stations, the scuba dive team found and collected krill from the underside of the ice.


3 September 2002

The LMG is occupying a relatively small floe (40 meter diameter) for a second consecutive day.  The scuba divers have had good success at this site and the ice deployed their third and final drifter buoy.  We were treated to a gorgeous sunrise (now about 07:30 LT) and a visit from a single emperor penguin in the morning.


4 September 2002

Somewhat surprisingly, the LMG has continuously occupied the same relatively small floe for a third consecutive day.  We have seen the pack loosen up and begin to re-freeze as the light winds have shifted slightly and the air temperatures have dropped.  This area looks completely different than it did just a short while ago when we arrived early this week.  The Ice team and the Krill team have had very good success here and the Seal team is pitching in to help wherever possible.


5 September 2002

The LMG moved off the floe we've occupied for the last three days, conducted CTD casts last evening and this morning, then repositioned alongside a large, thick floe with heavy snow cover to work for the day.  Three crabeater seals have hauled out just aft of the ship.  Meanwhile, the scuba team and the ice corers have been busy collecting samples.  The Seal Team was finally able to work on another crabeater in the evening.  Although all their satellite tags were deployed early in the cruise, they are still able to make physiological studies. We are seeing some rather significant relative motion between floes due to a slight swell running through the area.


6 September 2002

The LMG continues to operate in the same general area, though on a different floe from yesterday.  The Ice Team recruited help from others on a beautifully clear day and collected many meters of core, and the divers completed two dives.  Everyone is pleased with the amount of work we're accomplishing, but we're starting to think about port calls and heading home.


7 September 2002

The LMG found a good ice floe yesterday so we are continuing to occupy it today.  The Ice team is collecting cores and optics measurements, the Dive team currently made 3 dives, and the Seal team was able to complete measurements on another crabeater seal.  We are definitely making productive use of the last few science days of GLOBEC IV.


8 September 2002

The LMG was not able to maintain position against the floe we worked on the past two days, but we quickly found another location in the immediate vicinity this morning.  After lunch the Ice Team went out to make optics measurements and the Dive team made two dives.  We are seeing lots of seals in the area, and one hauled out relatively close to the ship so the Seal team was able to complete measurements on their 15th seal.


9 September 2002

We have completing our final day of on-ice science from the LMG, though analysis of samples in the lab will continue for some time yet.  The Dive team completed three dives and the Seal team completed their 16th and “last” seal.  We are communicating with both the NBP and Palmer Station to coordinate activities for the next few days, and to think about preparations for the upcoming port call in Punta Arenas.


Individual group reports: 


BG-232 – Seal Foraging Ecology (Costa, Shaffer, Barnes, McDonald, and Kuhn)

The last seal we worked with was on 16 August, since that time we have either not seen seals or were not in a position that allowed us to work with them.  This all changed when we came to Process Station 3.  In this region we have seen many seals in the water with a few hauled out that were too far away to safely reach.  Finally on 5 September a seal hauled out immediately behind the ship.  However, the ice was very loose and although the floe was behind the ship it would have required crossing several loose floes to reach it. After some discussion we decided to wait until the other research teams were off the ice or out of the water.  Once the other groups were back on board we were able to maneuver the ship to use the personnel basket to reach the flow. We were thus able to capture out first seal in almost 3 weeks.  Our luck continued as seals hauled out nearby the ship on 7 & 8 September.  Although we had deployed all of our tags we still had physiological measurements to complete.  These included measurements of the animals oxygen stores, which are the primary determinate of diving ability.  Measurements included blood volume, muscle myoglobin and blood hemoglobin content.  We used the time between captures to analyze our samples from the first part of the cruise allowing the development of other sampling protocols that we were able to test on these last three animals. We also collected data on the animals’ body mass and composition, as well as blubber biopsies for estimates of diet from free fatty acid analysis.  This will bring our total number of seals captured during this cruise to 16; twice the 8 we captured during the winter cruise last year.  Our total for the entire 2-year, 4 cruise GLOBEC field program is 44 crabeater seals and one leopard seal.  We are now cleaning up and packing our gear and looking forward to the trip home.


BG-234 – Penguin Foraging Ecology (Heidi Geisz, Brett Pickering)

The penguin team was able to finish deploying their transmitters this week while on the NBP.  During the afternoon of the 3rd a couple of small groups of Adélie Penguins were spotted, but it was decided to continue on in hopes of a larger group.  A group of 13 birds were spotted and the ship maneuvered into position so that the team assembled could be deployed onto the ice using the ship’s bow crane.  With seven people on the ice 9 Adélies were captured.  Four of the birds had transmitters affixed to their back feathers and the remaining 5 birds were diet sampled by the NBP bird team.  We would like to thank all the science groups of the NBP for having us aboard, fitting us into their tight schedule and helping us with our work.   We would like to especially thank the NBP bird team of Dr. Chris Ribic and Erik Chapman, who in addition to helping us during our week aboard also have taken diet samples for our group during all four GLOBEC cruises.  The Captain and crew’s efforts have also been much appreciated during our work on the NBP.



BG-235 - Sea Ice Microbial Communities (Adkins, Blees, Cunningham, Hartsough, Fritsen)

As we have drifted north, the floes have become smaller and the pack is generally breaking up.  It has become difficult to stay on one station.  We have taken on the strategy of quick (one day) ice stations, assuming that we will not be there long.  In the last week we have had 4 ice stations.  Ice sampling has continued although we have largely had to forgo transects and other physical characterization in favor of quickly isolating different biotic environments and sampling them.  Forty-eight ice cores were collected in efforts to determine integrated biomass of sea ice microalgal communities, and nutrients and salinity characterization.  Work continues in conjunction with OG-241 at the optics sites to co-characterize optical environments and their associated biota. In addition we continue work with BG-244 to try to sample from the surface algal communities that they have identified during dives.

We have continued to find enhanced algal biomass (>20 micro grams per liter) in cores measuring over 70 centimeters in length.  The distributions of the peaks in biomass in all cores have been distributed at all depths in the ice core profiles.  Production assays on these communities indicate high potential for in situ production and in situ photo-adaptation.  Further observations and experimentation at the northern most stations has noted the early onset of the spring bloom occurring in the ice.



OG-241 Optical Environment of the Western Antarctic Peninsula Region (Elder and Claffey)

In the last week we have concluded on ice operations for Southern GLOBEC.  We set up operations between grid stations 4 and 5.  In the last week we have operated in a mobile operation spending 1-2 days on a floe then moving on. The ice in this area experienced a swell for most of the week, waning mid week, then picking up again for the weekend.  The first 3 days we spent on a single ice floe where we deployed our last long term drifting ice buoy, performed optics, drilled thickness transects and sampled the ice for salinity, oxygen isotopes, structure, temperature profile, and other measurements performed by BG235, BG244. We then moved off to another floe that where we drilled 50 holed for ice thickness, we were assisted by BG232 (the seal team). This floe was the thickest that we have drilled on this cruise. Snow depth along the lines averaged 82cm (33 inches) of snow, with the greatest snow depth being 140cm (55 inches). Ice thickness along the lines averaged 2.59m (8.5 feet), with the greatest ice thickness measured being 5.65m (18.5 feet).  These thick floes are not all that common in this area (less than 10 percent) but seem to frequently be seen. Ice cores of the top 4 meters were also taken. This ice is older and must have drifted up from farther south. The floes have steep vertical side that indicate that they were once part of an ice sheet, not just a ridged and snow blown block of ice.  The next 2 days were spent at a single ice floe where more than a dozen ice cores were taken as well as an optics station.  The last 2 days of science on the ice were toned down a bit as we obtained more optics and less coring.  Samples were also taken this week that will be shipped home for further analysis. Here at the end of this week, we have drifted north of the process grid and the ice has loosened up greatly with bands of loose ice blocks and the swell has returned indicating that the open ocean cannot be too far away.  For the next week we will finish up sample work in the freezer, continue hourly ice observations until we travel out of the ice pack, back up data sets and make sure that everyone involved with co-located samples gets the data that they need for their project before getting back to land.


BG-244 - Winter Ecology of Larval Krill: Quantifying their interaction with the Pack Ice habitat (Quetin, Hessell, Willis, Boch, Oakes, Johnston and Dovel)

This week we occupied the northern-most Process Station, #3, and, for the first time on the cruise, found krill larvae directly associated with and feeding on the pack ice surface.  The weather during the week was relatively mild, and we dove daily.  We completed 18 dives this week and found larvae on every dive except one, which made for a very productive week.  During this station we collected krill for experiments on instantaneous growth rates, fecal pellet production and whole body fluorescence, preserved krill for analysis of condition factor and measured samples for length frequency and stage analysis.  Due to the relatively loose pack ice, this northern-most Process Station was a series of locations close to each other, so some of the transect methods used at other stations were not feasible.   At this station we were able to collect a set of 77 suction samples from ice surfaces, which will enable us to assess the variability of the microbial community within and between the three process stations occupied during the cruise.  Appropriately, on our final day of sampling, September 9, we were again swimming with a group of curious crabeater seals, a fitting end to a productive cruise with a great group of people on board.  To thank any one person for the success of this cruise would be inappropriate.  Thank you all, from the bridge to the mid-decks to the engine room for your help, good nature and willingness to put up with our crazy, hard working group.



Submitted by:

Daniel Costa, Chief Scientist and

Karl Newyear, Marine Projects Coordinator

LMG 02-05